OpenStudy (anonymous):

Determine the domain for f(x) given, f(x) = √x+1

jimthompson5910 (jim_thompson5910):

is the equation \[\large f(x) = \sqrt{x} + 1\] or is it \[\large f(x) = \sqrt{x + 1}\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

A square root must always have a positive value beneath it, otherwise your answer will contain an imaginary number. So, \[x+1 \ge 0 \] By subtracting the 1 from both sides, we find our domain, \[x \ge -1\]

OpenStudy (anonymous):

alright, thanks so much!!

OpenStudy (anonymous):

No problem, glad to help.